# On the Source Model Key Agreement Problem

### v2

@@ Abstract
We consider the source model key agreement problem involving two legitimate parties and an eavesdropper who observe $n$ i.i.d. samples of $X$ and $Y$ and $Z$ respectively. In this paper, we focus on one of the simplest instances where the key capacity remains open, specifically when $X$ and $Y$ are binary random variables and $Z$ is a function of the pair $(X,Y)$. The best-known upper bound on the key capacity is characterized by an inf-max optimization problem that generally lacks a closed-form solution. We provide general conditions under which the upper bound reduces to $I(X;Y)$. As an example, we consider the XOR setting in which $X$ and $Y$ are binary, and $Z$ is the XOR of $X$ and $Y$. The upper bound reduces to $I(X;Y)$ for this source. Next, we conjecture that the rate $I(X;Y)$ is not achievable for the XOR source, and provide some ideas that might be useful for developing a new upper bound on the source model problem.
@@ 1 Introduction
The two-party source model problem consists of two legitimate parties, Alice and Bob, who aim to establish a shared secret key hidden from an eavesdropper, Eve [1, 2]. Alice, Bob, and Eve observe $n$ i.i.d. samples of $X$ and $Y$ and $Z$ respectively. Random variables $X,Y$ and $Z$ are distributed according to $P_{X,Y,Z}$, known to all parties. Through authenticated but public communication, Alice and Bob collaborate to generate a key that remains secret from Eve (extensions to multiple parties and continuous random variables can be found in [3, 4, 5, 6, 7, 8]; please see [9] for a detailed treatment). In a non-interactive setting, Alice and Bob locally generate messages $F_{1}$ and $F_{2}$ from $P_{F_{1}|X^{n}}$ and $P_{F_{2}|Y^{n}}$, respectively, and exchange these messages. Then Alice and Bob create their own key from $K_{A}\sim P_{K_{A}|X^{n},F_{1},F_{2}}$ and $K_{B}\sim P_{K_{B}|Y^{n},F_{1},F_{2}}$ respectively. On the other hand, in the interactive setting, Alice initiates the process by generating a public message $F_{1}$ using the distribution $P_{F_{1}|X^{n}}$ and shares it with Bob. Bob then responds with $F_{2}$ derived from $P_{F_{2}|Y^{n},F_{1}}$. This exchange continues for $r$ rounds, during which both senders alternately generate and share public messages $F_{3},F_{4},\cdots F_{r}$. At the end of these interactions, Alice derives a key $K_{A}$ based on $P_{K_{A}|X^{n},F_{1:r}}$ while Bob produces a corresponding key $K_{B}$ using $P_{K_{B}|Y^{n},F_{1:r}}$. In an $(n,R,\delta)$ code, the keys must match with probability $1-\delta$:
@@
where $\delta$ is a small positive value. The key rate is defined as $R=\frac{1}{n}H(K_{A})$. $R$ must exhibit high entropy, quantified as:
@@
Note that the value of $r$ for non-interactive regime is $2$. We note that if we omit the factor $1/n$ in equation (3), we obtain a seemingly stronger notion of secrecy; however, this is equivalent to the weaker form presented in (3). For a proof of equivalence, see [10].
@@
Finding the secret key capacity remains an open problem in general scenarios, including both the interactive and non-interactive communication settings. Conditions that imply $S(X;Y\|Z)>0$ are studied in [2, 11, 12, 13]. The exact formula for secret key capacity is known only for some instances, such as when $X,Y$ and $Z$ form a Markov chain in some order. As a side result, we derive the capacity for a new class of sources, which we call “deterministic-erasure” sources in Appendix A.
@@
where the conditions $|\mathcal{V}|\leq|\mathcal{X}|$ and $|\mathcal{U}|\leq|\mathcal{X}|$ can be imposed on $\mathcal{V}$, $\mathcal{U}$, and $\mathcal{X}$ representing the alphabets associated with $V$, $U$, and $X$, respectively [9]. [^1] We can equivalently express $S_{\text{ow}}(X\rightarrow Y\|Z)=\max_{V-\!\!\!\!\minuso\!\!\!\!-U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}I(U;Y|V)-I(U;Z|V),$ (5) This follows from noting that changing $U$ to $(U,V)$ will not change the mutual information terms, while introducing the additional Markov chain constraint. However, this comes at the cost of a larger cardinality bound on $U$. It is evident that $S_{\text{ow}}(X\rightarrow Y\|Z)$ establishes a lower bound on the SK capacity $S(X;Y\|Z)$.
@@
$\min_{J-\!\!\!\!\minuso\!\!\!\!-Z-\!\!\!\!\minuso\!\!\!\!-(X,Y)}I(X;Y|J)$
is called the intrinsic conditional mutual information. More generally, we can relax the degraded assumption $J-\!\!\!\!\minuso\!\!\!\!-Z-\!\!\!\!\minuso\!\!\!\!-(X,Y)$ to a less noisy condition. The inequality
@@
holds if the channel $P_{Z|X,Y}$ is less noisy than the channel $P_{J|X,Y}$. Even more generally, for any arbitrary $P_{J|X,Y}$, we have [14]
$S(X;Y\|Z)-S(X;Y\|J)\leq\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)}I(U;J|V)-I(U;Z|V)=S_{\text{ow}}(XY\rightarrow Z|J)$ (6)
Observe that the right hand side of the above equation vanishes when $P_{Z|X,Y}$ is less noisy than the channel $P_{J|X,Y}$, and was interpreted as “the penalty of deviating from the less-noisy condition” in [13]. Using (6) and the upper bound
$S(X;Y|J)\leq I(X;Y|J)$
the following upper bound was proposed in [14]:
$S(X;Y|Z)\leq\Psi(X;Y\|Z)$ $\triangleq\inf_{J}\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)}I(X;Y|J)+I(U;J|V)-I(U;Z|V)$ (7)
Note that $J$ serves as an auxiliary receiver in the sense of [15]. The above bound is the best known upper on the key capacity. It involves a min-max optimization problem, which makes computing the upper bound challenging. In fact, it is not known whether the infimum over $J$ in (7) is a minimum (see [14] for a discussion on the computability of this bound). A weaker, but more convenient upper bound was also proposed in [14]
@@
Notice that the upper bound $\Psi(X;Y\|Z)$ in (7) depends solely on $P_{X,Y,J}$ and $P_{X,Y,Z}$, meaning the minimization is restricted to $P_{J|X,Y}$. However the weaker upper bound $\hat{\Psi}(X;Y\|Z)$ in (9) requires minimization over $P_{J|X,Y,Z}$. On the other hand, the weaker upper bound $\hat{\Psi}(X;Y\|Z)$ does not involve the auxiliary random variables $U$ and $V$.
The entire research presented in this paper is motivated by our desire to study $S(X;Y\|Z)$ in what we believe is the simplest case where the key capacity remains unknown, namely, when $X$ and $Y$ are binary random variables, and $Z$ is a function of the pair $(X,Y)$. The main results of this paper are as follows:

### v3

@@ Abstract
We consider the source model key agreement problem involving two legitimate parties and an eavesdropper who observes $n$ i.i.d. samples of $X$ and $Y$ and $Z$, respectively. In this paper, we focus on one of the simplest instances where the key capacity remains open, specifically when $X$ and $Y$ are binary random variables and $Z$ is a function of the pair $(X,Y)$. The best-known upper bound on the key capacity is characterized by an inf-max optimization problem that generally lacks a closed-form solution. We provide general conditions under which the upper bound reduces to $I(X;Y)$. As an example, we consider the XOR setting in which $X$ and $Y$ are binary, and $Z$ is the XOR of $X$ and $Y$. The upper bound reduces to $I(X;Y)$ for this source. Next, we conjecture that the rate $I(X;Y)$ is not achievable for the XOR source, and provide some ideas that might be useful for developing a new upper bound on the source model problem.
@@ 1 Introduction
The two-party source model problem consists of two legitimate parties, Alice and Bob, who aim to establish a shared secret key hidden from an eavesdropper, Eve [1, 2]. Alice, Bob, and Eve observe $n$ i.i.d. samples of $X$ and $Y$ and $Z$, respectively. Random variables $X,Y$, and $Z$ are distributed according to $P_{X,Y,Z}$, known to all parties. Through authenticated but public communication, Alice and Bob collaborate to generate a key that remains secret from Eve (extensions to multiple parties and continuous random variables can be found in [3, 4, 5, 6, 7, 8]; please see [9] for a detailed treatment). In a non-interactive setting, Alice and Bob locally generate messages $F_{1}$ and $F_{2}$ from $P_{F_{1}|X^{n}}$ and $P_{F_{2}|Y^{n}}$, respectively, and exchange these messages. Then Alice and Bob create their own key from $K_{A}\sim P_{K_{A}|X^{n},F_{1},F_{2}}$ and $K_{B}\sim P_{K_{B}|Y^{n},F_{1},F_{2}}$, respectively. On the other hand, in the interactive setting, Alice initiates the process by generating a public message $F_{1}$ using the distribution $P_{F_{1}|X^{n}}$ and shares it with Bob. Bob then responds with $F_{2}$ derived from $P_{F_{2}|Y^{n},F_{1}}$. This exchange continues for $r$ rounds, during which both senders alternately generate and share public messages $F_{3},F_{4},\cdots,F_{r}$. At the end of these interactions, Alice derives a key $K_{A}$ based on $P_{K_{A}|X^{n},F_{1:r}}$, while Bob produces a corresponding key $K_{B}$ using $P_{K_{B}|Y^{n},F_{1:r}}$. In an $(n,R,\delta)$ code, the keys must match with probability $1-\delta$:
@@
where $\delta$ is a small positive value. The key rate is defined as $\frac{1}{n}H(K_{A}). RR$ must exhibit high entropy, quantified as:
@@
Note that the value of $r$ for the non-interactive regime is $2$. We note that if we omit the factor $1/n$ in equation (3), we obtain a seemingly stronger notion of secrecy; however, this is equivalent to the weaker form presented in (3). For a proof of equivalence, see [10].
@@
Originally studied by Ahlswede and Csiszár [1] and Maurer [2], determining the secret key capacity remains a fundamental open problem in network information theory. For an introduction to this problem, the reader may refer to Chapter 22 of [9]. In general scenarios, the key capacity remains unknown (both in the interactive and non-interactive communication settings) with the exception of certain special cases, such as when $X,Y,$ and $Z$ form a Markov chain in some order. As a side result, we derive the capacity for a new class of sources, which we call “deterministic-erasure” sources in Appendix A.
@@
where the conditions $|\mathcal{V}|\leq|\mathcal{X}|$ and $|\mathcal{U}|\leq|\mathcal{X}|$ can be imposed on $\mathcal{V}$, $\mathcal{U}$, and $\mathcal{X}$ representing the alphabets associated with $V$, $U$, and $X$, respectively [9]. [^1] We can equivalently express $S_{\text{ow}}(X\rightarrow Y\|Z)=\max_{V-\!\!\!\!\minuso\!\!\!\!-U-\!\!\!\!\minuso\!\!\!\!-X-\!\!\!\!\minuso\!\!\!\!-(Y,Z)}I(U;Y|V)-I(U;Z|V),$ (5) This follows from noting that changing $U$ to $(U,V)$ will not change the mutual information terms, while introducing the additional Markov chain constraint. However, this comes at the cost of a larger cardinality bound on $U$. It is evident that $S_{\text{ow}}(X\rightarrow Y\|Z)$ establishes a lower bound on the SK capacity $S(X;Y\|Z)$. In particular, setting $U=X$ and $V$ constant yields
$S(X;Y\|Z)\geq I(X;Y)-I(X;Z).$
The best known lower bound is obtained in [11] as follows: consider random variables $U_{1},U_{2},\ldots,U_{k}$ satisfying the Markov chain conditions
$U_{i}-\!\!\!\!\minuso\!\!\!\!-(X,U_{1:i-1})-\!\!\!\!\minuso\!\!\!\!-(Y,Z)\quad\text{for odd }i,$ (6)
$U_{i}-\!\!\!\!\minuso\!\!\!\!-(Y,U_{1:i-1})-\!\!\!\!\minuso\!\!\!\!-(X,Z)\quad\text{for even }i.$ (7)
For any integers $1\leq m\leq k$, the secret key capacity satisfies $S(X;Y\|Z)$ is bounded from below by
$\Bigg{[}\sum_{\begin{subarray}{c}i\geq m\\ \text{odd }i\end{subarray}}I(U_{i};Y|U_{1:i-1})-I(U_{i};Z|U_{1:i-1})\Bigg{]}+\Bigg{[}\sum_{\begin{subarray}{c}i\geq m\\ \text{even }i\end{subarray}}I(U_{i};X|U_{1:i-1})-I(U_{i};Z|U_{1:i-1})\Bigg{]}.$ (8)
This bound is difficult to evaluate since both $m$ and $k$ are arbitrary and evaluation of the bound becomes increasingly difficult for large $m,k$.
@@
$I(X;Y\downarrow Z)\triangleq\min_{J-\!\!\!\!\minuso\!\!\!\!-Z-\!\!\!\!\minuso\!\!\!\!-(X,Y)}I(X;Y|J)$ (9)
is called the intrinsic information. More generally, we can relax the degraded assumption $J-\!\!\!\!\minuso\!\!\!\!-Z-\!\!\!\!\minuso\!\!\!\!-(X,Y)$ to a less noisy condition. The inequality
@@
holds if the channel $P_{Z|X,Y}$ is less noisy than the channel $P_{J|X,Y}$. Even more generally, for any arbitrary $P_{J|X,Y}$, we have [11, Corollary 2]
$S(X;Y\|Z)-S(X;Y\|J)\leq\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)}I(U;J|V)-I(U;Z|V)=S_{\text{ow}}(X,Y\rightarrow Z|J)$ (10)
Observe that the right-hand side of the above equation vanishes when $P_{Z|X,Y}$ is less noisy than the channel $P_{J|X,Y}$, and was interpreted as “the penalty of deviating from the less-noisy condition” in [13]. Using (10) and the upper bound
$S(X;Y\|J)\leq I(X;Y|J)$
the following upper bound was proposed in [11, Corollary 2]:
$S(X;Y\|Z)\leq\Psi(X;Y\|Z)$ $\triangleq\inf_{J}\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)}I(X;Y|J)+I(U;J|V)-I(U;Z|V)$ (11)
Note that $J$ serves as an auxiliary receiver in the sense of [14]. The above bound is the best known upper bound on the key capacity. It involves a min-max optimization problem, which makes computing the upper bound challenging. In fact, it is not known whether the infimum over $J$ in (11) is a minimum (see [11] for a discussion on the computability of this bound). A weaker, but more convenient upper bound was also proposed in [11]
@@
Notice that the upper bound $\Psi(X;Y\|Z)$ in (11) depends solely on $P_{X,Y,J}$ and $P_{X,Y,Z}$, meaning the minimization is restricted to $P_{J|X,Y}$. However, the weaker upper bound $\hat{\Psi}(X;Y\|Z)$ in (13) requires minimization over $P_{J|X,Y,Z}$. On the other hand, the weaker upper bound $\hat{\Psi}(X;Y\|Z)$ does not involve the auxiliary random variables $U$ and $V$.
Remark 1.
There are notable parallels between key agreement in classical and quantum scenarios [15, 16]. Furthermore, the quantum counterparts of the upper bounds in the key agreement problem represent independently interesting objects of study; while intrinsic information serves primarily as an upper bound on $S(X;Y\|Z)$ in the classical setting, its quantum counterpart—squashed entanglement—acts as a meaningful measure of entanglement for bipartite quantum systems [17, Section III]. On the one hand, the intrinsic information upper bound has led to the definition of squashed entanglement. On the other hand, insights from quantum information theory have been used to derive new upper bounds for the classical key capacity $S(X;Y\|Z)$, as demonstrated in [18, 19]. In [20], it is demonstrated that in the classical framework, if we permit $J$ to be a quantum system while maintaining $(X,Y,Z)$ as classical, the upper bound on $S(X;Y\|Z)$ in equation (14) remains valid. However, it is important to note that the minimum in (14) should be replaced by an infimum, as there is currently no known dimension bound on the size of the quantum system $J$ (see [21]).
Remark 2.
The question of when key agreement is possible, i.e., when $S(X;Y\|Z)>0$, has been an important topic in the literature [2, 12, 22, 13]. Orlitsky and Wigderson show that if $S(X;Y\|Z)>0$, just two rounds of interactive communication suffice to achieve a positive key rate [23]. Although multi-letter characterizations exist for the condition $S(X;Y\|Z)=0$ [13, 23], these results are not in single-letter form and thus cannot be used to verify whether $S(X;Y\|Z)=0$. The only known method to prove $S(X;Y\|Z)=0$ is to show that one of the known upper bounds on $S(X;Y\|Z)$ vanishes. It is indeed conjectured in [12] that $S(X;Y\|Z)=0$ if and only if the intrinsic information vanishes, i.e., $I(X;Y\downarrow Z)=0$. Given the upper bound in Proposition 1, the condition $I(X;Y\downarrow Z)=0$ in the conjecture should be refined to state that there exists a random variable $J$ such that (i) $X\rightarrow J\rightarrow Y$ forms a Markov chain, and (ii) the channel $P_{Z|X,Y}$ is less noisy than the channel $P_{J|X,Y}$.
1.1 Overview of our contributions
Consider the upper bound
$S(X;Y\|Z)\leq\Psi(X;Y\|Z)=\inf_{J}\;\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)}\Big{\{}I(X;Y|J)+I(U;J|V)-I(U;Z|V)\Big{\}}.$
No cardinality bound on the alphabet of $J$ is known. However, even after fixing a choice of $P_{J|X,Y,Z}$, the cardinality of the alphabets of $U$ and $V$ in the inner maximization can be bounded from above by $|\mathcal{X}||\mathcal{Y}|$. Thus, after fixing $J$, one still needs to solve an optimization over $P_{U,V|X,Y,Z}$ with $|\mathcal{X}||\mathcal{Y}||\mathcal{Z}|\cdot(|\mathcal{X}||\mathcal{Y}|)^{2}$ free parameters, rendering a direct evaluation of the upper bound practically infeasible.
Our first observation is that when $Z$ is a function of $(X,Y)$, the above upper bound coincides with the simpler upper bound
$S(X;Y\|Z)\leq\hat{\Psi}(X;Y\|Z)=\min_{J}\;I(X;Y|J)+I(X,Y;J|Z),$
which eliminates the auxiliary variables $(U,V)$ in the inner maximization. When $|\mathcal{X}|$ and $|\mathcal{Y}|$ are small and $Z$ is a function of $(X,Y)$, it is feasible to numerically evaluate $\hat{\Psi}(X;Y\|Z)$; in this case one can impose a cardinality bound $|\mathcal{J}|\leq|\mathcal{X}||\mathcal{Y}|$ on $J$. However, $I(X;Y|J)+I(X,Y;J|Z)$ is neither convex nor concave in $P_{J|X,Y,Z}$. Consequently, even if one has a good guess for the minimizing $P_{J|X,Y,Z}$, it may be hard to certify it as a global minimizer.
Our first contribution is a “weak duality” technique to address this certification problem. We believe the technique is of independent interest and may be useful for other non-convex optimization problems in network information theory.
1.1.1 Main Contribution 1
Many problems in network information theory involve rate regions or non-convex optimizations over auxiliary random variables. As an example, consider the broadcast channel: Marton’s inner bound involves three auxiliary random variables representing private and common messages. For certain structured channels (e.g., semi-deterministic or erasure Blackwell channels), the global optimizers take a particular form. In some cases, this structure is provable; in others, it is only supported by numerical evidence. To address questions of this type, we propose a “weak-duality” idea and illustrate it in the source model.
Statement of the issue: Let $\mathcal{P}$ denote the set of joint distributions $P_{X,Y,Z}$ for which the global minimizer of $\hat{\Psi}(X;Y\|Z)$ is a constant $J$. Our goal is to characterize the set $\mathcal{P}$. Equivalently, we seek to identify all $P_{X,Y,Z}$ satisfying
$I(X;Y)-I(X;Y|J)-I(X,Y;J|Z)\leq 0,\quad\forall\,P_{J|X,Y,Z}.$
The equality in this inequality holds when $J$ is a constant. Hence, we may rewrite the condition as
$I(X;Y)-H(X,Y|Z)=\min_{J}\big{[}I(X;Y|J)+H(X,Y|J,Z)\big{]}.$
Observe that $\min_{J}\big{[}I(X;Y|J)-H(X,Y|J,Z)\big{]}$ computes the lower convex envelope of the mapping
$Q_{X,Y,Z}\mapsto I_{Q}(X;Y)-H_{Q}(X,Y|Z)$ (15)
evaluated at $Q_{X,Y,Z}=P_{X,Y,Z}$ (see [24] for details on the relation between lower convex envelopes and auxiliary random variables). The mapping in (15) is defined for all $Q_{X,Y,Z}$ on fixed alphabets $\mathcal{X}\times\mathcal{Y}\times\mathcal{Z}$.
Given a function $\varphi(\cdot)$, its lower convex envelope $\mathcal{K}[\varphi]$ is defined as the largest convex function dominated by $\varphi$, i.e.,
$\mathcal{K}[\varphi](x)=\sup\{\tilde{\varphi}(x)\mid\tilde{\varphi}\text{ convex and }\tilde{\varphi}\leq\varphi\}.$
See Figure 1 for an illustration.
The set $\mathcal{P}$ consists precisely of those distributions $P_{X,Y,Z}$ where the mapping in (15) coincides with its lower convex envelope. For a general non-convex function $\varphi(x)$, computing $\mathcal{K}[\varphi](x^{\star})$ at a point $x^{\star}$ requires global knowledge of $\varphi(\cdot)$; local information about its value near $x^{\star}$ is insufficient. Thus, determining the set $\{x:\mathcal{K}[\varphi](x)=\varphi(x)\}$ necessitates a global access to the function $\varphi(\cdot)$. However, given that the mapping in (15) is not arbitrary but constructed from Shannon entropy measures, can this structural property enable us to derive global conclusions from local analysis?
Figure 1: Visualization of a function and its lower convex envelope. The function coincides with its lower convex envelope for $x\geq 1$ and $x\leq-1$.
From local to global: Our approach to bridge global to local is as follows. Define
$f(X,Y,Z,J)\triangleq I(X;Y)-I(X;Y|J)-I(X,Y;J|Z).$
We seek the set of $P_{X,Y,Z}$ such that
$\max_{J}f(X,Y,Z,J)\;\leq\;0,$
where the maximization is over $P_{J|X,Y,Z}$. Suppose we can find another functional $g(X,Y,Z,T)$ such that for every $P_{X,Y,Z}$ we have
$\max_{J}f(X,Y,Z,J)\;\leq\;\min_{T}g(X,Y,Z,T),$
where the minimization is over $P_{T|X,Y,Z}$. This is akin in spirit to linear-programming duality, but carried out at the level of mutual-information expressions. If such a $g$ exists, then for a given $P^{\star}_{X,Y,Z}$ one can certify
$\max_{J}f(X,Y,Z,J)\;\leq\;0$
by presenting a single witness $P_{T|X,Y,Z}$ for which
$g(X,Y,Z,T)=0$
holds at $P^{\star}_{X,Y,Z}$. Crucially, only a single witness and a local verification at $P^{\star}_{X,Y,Z}$ is needed.
In this paper, we take
$g(X,Y,Z,T)\triangleq I(T;Z)+I(X;Y|T)+I(T;Z|X,Y).$
One then needs to verify that for every $P_{X,Y,Z,T,J}$,
$I(X;Y)-I(X;Y|J)-I(X,Y;J|Z)\;\leq\;I(T;Z)+I(X;Y|T)+I(T;Z|X,Y).$
The above is a non-Shannon-type inequality. It can, however, be established using Shannon-type inequalities if we impose the constraint $I(T;J|X,Y,Z)=0$. This constraint can be added without loss of generality because the expression depends only on the marginals $P_{X,Y,Z,J}$ and $P_{X,Y,Z,T}$. See the proof of Theorem 1.
To demonstrate the power of this duality idea, we consider the case where $X$ and $Y$ are binary and $Z$ is a function of $(X,Y)$. We show that a tight characterization of $\mathcal{P}$ can be obtained in the binary case. Namely, if $P_{X,Y,Z}\in\mathcal{P}$, i.e., constant $J$ is optimal, we can explicitly identify a witness $P_{T|X,Y,Z}$ for which $g(X,Y,Z,T)=0$. In addition, we also study cases where $P_{X,Y,Z}\notin\mathcal{P}$, i.e., a non-constant $J$ is the minimizer; here, we find the minimizer $J$ as well and show that the minimizer makes $I(X;Y|J)+I(X,Y;J|Z)=0$. Thus, we have $S(X;Y\|Z)=0$ for $P_{X,Y,Z}\notin\mathcal{P}$. See Theorem 2 for details. A particular example of interest is when $Z=X\oplus Y$. The set $\mathcal{P}$ restricted to $Z=X\oplus Y$ is the entire set of all joint distributions on $\mathcal{X}\times\mathcal{Y}$ and we obtain $\hat{\Psi}(X;Y\|Z)=\Psi(X;Y\|Z)=I(X;Y)$ for every $P_{X,Y}$, when $Z=X\oplus Y$.
1.1.2 Main Contribution 2
Having established that the best known upper bound on $S(X;Y\|Z)$ reduces to $I(X;Y)$ in some cases, our second contribution is evidence of the suboptimality of this upper bound. We also provide a new upper bound for the non-interactive case, which is always no larger than the best known bound. Unfortunately, evaluating our new bound is challenging because it has essentially the same complexity as $\Psi(X;Y\|Z)$, which, as discussed earlier, is computationally intractable.
As discussed earlier, we use the “weak-duality” idea to show that when $X$ and $Y$ are binary and $Z=X\oplus Y$, we have $\hat{\Psi}(X;Y\|Z)=\Psi(X;Y\|Z)=I(X;Y)$. Can it be that $I(X;Y)$ is also achievable, i.e., $S(X;Y\|Z)=I(X;Y)$?
Achievability of the key rate $I(X;Y)$. If $Z$ is a constant random variable, then $S(X;Y\|Z)=I(X;Y)$. One strategy to achieve this rate has one party (say Alice) randomly bin $X^{n}$ at the Slepian–Wolf rate of $nH(X|Y)$ bits and communicate the bin index to Bob over the public channel. Bob then recovers $X^{n}$ and the parties generate $nH(X)$ bits of common randomness. Since $nH(X|Y)$ bits are revealed publicly, they perform privacy amplification and output n(H(X)H(X|Y))=nI(X;Y)n(H(X)-H(X|Y))=nI(X;Y) secret bits that remain hidden from the eavesdropper. This protocol is communication-heavy: to generate $nI(X;Y)$ secret bits, it reveals $nH(X|Y)$ bits about $X^{n}$ on the public channel. A symmetric protocol with Bob speaking reveals $nH(Y|X)$ bits. One may ask whether interactive communication can reduce the public communication needed to achieve $I(X;Y)$. While this is possible in certain instances, the answer is negative in general: the minimum public communication rate to achieve $I(X;Y)$ is characterized in [6]. It has been conjectured in [6] (and proved in some cases) that when $X$ and $Y$ are binary, at least $n\min\{H(X|Y),H(Y|X)\}$ bits must be revealed publicly to achieve a key rate of $I(X;Y)$.
Now suppose $Z$ is a general, non-constant random variable (i.e., the eavesdropper has side information $Z$ about $(X,Y)$ in addition to access to the public transcript). Clearly, $S(X;Y\|Z)\leq I(X;Y)$. But can we still achieve $I(X;Y)$? This specific question does not appear to have been explicitly addressed in the literature. We conjecture that when $X$ and $Y$ are binary and $Z=X\oplus Y$, the rate $I(X;Y)$ is not achievable for a general $P_{X,Y}$. To begin, consider the one-way protocols above when $Z=X\oplus Y$. If $H(Z)>\max\{H(X),H(Y)\}$, then $H(X|Z)<H(X|Y)$ and $H(Y|Z)<H(Y|X)$. In this case, the one-way protocols fail because, assuming that the bin index of $X^{n}$ at rate $H(X|Y)$ is revealed, the eavesdropper can utilize $Z^{n}$ and a Slepian–Wolf decoder to reconstruct the entire sequence $X^{n}$.
Next, consider interactive protocols. One strategy is to design an interactive transcript $F_{1},F_{2},\ldots,F_{r}$ that yields secret keys such that
$\mathbb{P}[K_{A}=K_{B}]$ $\approx 1,$
$\frac{1}{n}H(K_{A})$ $\approx I(X;Y),$
$\frac{1}{n}I(K_{A};F_{1:r})$ $\approx 0,$
$\frac{1}{n}H(Z^{n}|F_{1:r})$ $\approx 0,$
so that the parties both (i) generate keys of maximum rate $I(X;Y)$ that are independent of the transcript $F_{1:r}$, and (ii) interactively compute the XOR $Z^{n}$ (a function-computation problem). If both goals are met, then $K_{A},K_{B}$ are independent of $(F_{1:r},Z^{n})$. But can these two goals be met simultaneously?
First, consider the non-interactive case with two messages $F_{1}$ and $F_{2}$ generated via $P_{F_{1}|X^{n}}$ and $P_{F_{2}|Y^{n}}$, respectively. The goal is to compute $Z^{n}=X^{n}\oplus Y^{n}$ from $(F_{1},F_{2})$ alone. This is the classical modulo-two sum problem in source coding, originally studied by Körner and Marton [25]. One can import ideas from this line of work (e.g., [26, 27]) to study the source model problem; this is one of our key ideas leading to a new upper bound in the non-interactive case.
For the interactive case with multiple rounds, we assume
$\frac{1}{n}H(Z^{n}|F_{1:r})\approx\Delta$
for some $\Delta\geq 0$. The case $\Delta=0$ corresponds to an interactive modulo-two sum problem. We provide evidence that $\Delta$ cannot be close to zero when achieving the rate $I(X;Y)$. Moreover, we derive a new upper bound for interactive protocols as a function of a given $\Delta$. To explain this bound, we first overview the existing best-known upper bound.
Overview of the proof of the best-known upper bound. Roughly speaking, the main steps to prove $S(X;Y\|Z)\leq\Psi(X;Y\|Z)$ are as follows: [^2] The proof sketched here is not the original proof given in [11]; instead, it follows the add-and-subtract “auxiliary receiver” approach; see, e.g., Section II.A in [28]. take an arbitrary $(n,R,\delta)$ code generating keys $K_{A}\approx K_{B}$ via interactive communication $F_{1:r}$. Then
$R\;\approx\;\frac{1}{n}H(K_{A})\;\approx\;\frac{1}{n}H(K_{A}|Z^{n},F_{1:r})$ $=\frac{1}{n}H(K_{A}|J^{n},F_{1:r})+\frac{1}{n}\big{(}H(K_{A}|Z^{n},F_{1:r})-H(K_{A}|J^{n},F_{1:r})\big{)}.$
The difference term can be single-letterized:
$\frac{1}{n}\big{(}H(K_{A}|Z^{n},F_{1:r})-H(K_{A}|J^{n},F_{1:r})\big{)}$ $=\frac{1}{n}\big{(}I(K_{A};J^{n}|F_{1:r})-I(K_{A};Z^{n}|F_{1:r})\big{)}$
$=\frac{1}{n}\sum_{i=1}^{n}\big{(}I(K_{A};J^{i},Z_{i+1}^{n}|F_{1:r})-I(K_{A};J^{i-1},Z_{i}^{n}|F_{1:r})\big{)}$
$=\frac{1}{n}\sum_{i=1}^{n}\big{(}I(K_{A};J_{i}|F_{1:r},J^{i-1},Z_{i+1}^{n})-I(K_{A};J^{i-1}Z_{i}|F_{1:r},J^{i-1},Z_{i+1}^{n})\big{)}$
$=I\big{(}K_{A};J_{Q}|Q,F_{1:r},J^{Q-1},Z_{Q+1}^{n}\big{)}-I\big{(}K_{A};Z_{Q}|Q,F_{1:r},J^{Q-1},Z_{Q+1}^{n}\big{)},$
where $Q$ is a time-sharing variable independent of $(K_{A},K_{B},F_{1:r},X^{n},Y^{n},Z^{n})$. Identifying $U=K_{A}$ and $V=(Q,F_{1:r},J^{Q-1},Z_{Q+1}^{n})$ yields
$\frac{1}{n}\big{(}H(K_{A}|Z^{n},F_{1:r})-H(K_{A}|J^{n},F_{1:r})\big{)}\;\leq\;\max_{(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)}\big{\{}I(U;J|V)-I(U;Z|V)\big{\}}.$
Next,
$\frac{1}{n}H(K_{A}|J^{n},F_{1:r})$ $\approx\frac{1}{n}I(K_{A};K_{B}|J^{n},F_{1:r})$
$\leq\frac{1}{n}I(X^{n};Y^{n}|J^{n},F_{1:r})$
$\leq\frac{1}{n}I(X^{n};Y^{n}|J^{n})$
$=I(X;Y|J),$
where the inequality $I(X^{n};Y^{n}|J^{n},F_{1:r})\leq I(X^{n};Y^{n}|J^{n})$ can be proved by induction on $r$.
Improvement to the upper bound. Our new idea is to exploit
$\frac{1}{n}H(Z^{n}|F_{1:r})\approx\Delta$
and impose a corresponding constraint on the auxiliary random variable $V=(Q,F_{1:r},J^{Q-1},Z_{Q+1}^{n})$. This is formalized in Theorem 6. We provide explicit upper bounds in the non-interactive case. Obtaining explicit bounds for the interactive case requires explicit bounds on $\Delta$ for protocols achieving the key rate $I(X;Y)$, which we leave for future work.
We also discuss other ideas for improving the upper bound. For instance, we show that using multi-letter versions of $\hat{\Psi}$ does not yield improved bounds. Moreover, as discussed above, the identity (10) is the key step in deriving the best-known upper bound: in that step we replace $Z$ with an auxiliary receiver $J$. To generalize this argument, one may also replace $X$ and $Y$ with other variables. We propose the following conjecture: for any $P_{X,Y,Z,X^{\prime},Y^{\prime},Z^{\prime}}$,
$S(X;Y\|Z)-S(X^{\prime};Y^{\prime}\|Z^{\prime})\;\leq\;S_{\text{ow}}(X,Y\!\to\!Z^{\prime}\|Z)+I(Y^{\prime},Z^{\prime};X|X^{\prime})+I(X^{\prime},Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime},Y^{\prime},Z^{\prime}).$
This conjecture generalizes (10) (it reduces to (10) by setting $X=X^{\prime}$ and $Y=Y^{\prime}$).
Summary of our contributions: The main results of this paper can be summarized as follows: